# Steering Generative Models with Experimental Data for Protein Fitness Optimization

### v1

@@ A.2 Protein fitness optimization task
We studied fitness optimization across two different protein-fitness datasets, TrpB and CreiLOV (Table 2). TrpB is 389 residues in length, but based on available fitness data, we limited design to 15 residues: 117, 118, 119, 162, 166, 182, 183, 184, 185, 186, 227, 228, 230, 231, and 301. Namely, we combined the fitness data from 6 combinatorially complete 3-site libraries (D-I from Johnston et al. (2024)) and the 4-site library across residues 183, 184, 227, and 228. We normalized the parent fitness to 1 in each dataset and rounded all negative fitness values up to zero. The fitness here is the catalytic rate of a native reaction, the formation of tryptophan from indole and serine. To obtain a proxy fitness for all variants in the design space ($20^{15}$ possibilities) we trained an oracle inspired by the dataset splitting and model architecture used in Blalock et al. (2025). Namely, we used all of the single, double, and triple mutants in the library for training, with 10% and 20% of the quadruple mutants being used for validation and testing, respectively. Our model consists of an ensemble of 20 MLPs for TrpB, and each was trained on one-hot encodings of the designed residues for 1000 epochs.
@@
Figure A1: Oracles trained on available labeled fitness data for TrpB and CreiLOV extrapolate well to higher order combinations of mutations within the design space, as measured by Pearson correlation.
@@
Our primary method for evaluation involved examining the distribution of sampled sequences and their corresponding fitness values, diversities, and novelties. The processing pipeline for generated sequences in shown in Fig. A2. In diffusion models, sequences were generated with fixed length equal to the parent length. For the language model, using the same starting sequence context in the ARLM (before the variable region) resulted in all generations mapping to the same (parent) sequence, so this strategy was not used. Instead, sequence fragments of variable length were generated and aligned with the parent sequence using mafft (Katoh & Standley, 2013). Then, gaps were replaced with the corresponding amino acid in the parent sequence to generate complete pseudo-sequences of a fixed length. Special tokens, which occurred rarely in generation, were replaced by a random amino acid. For TrpB, residues outside of the design space of 15 residues were naively mapped to the original amino acid type in the parent sequence at the end of generation. We did not test inpainting, although this could be accomplished with masked (diffusion) language models.
@@
There are several reasons why we did not directly compare the performance of SGPO methods to existing methods used in protein engineering, such as directed evolution and MLDE. In the case of directed evolution (such as random mutagenesis): (1) It is not obvious which parent sequences to use as the starting points for directed evolution for a fair comparison. (2) It is unclear if the oracle captures the true nature of the protein fitness landscape or extrapolates well to sequences with many mutations relative to the original fitness dataset from which the oracle was trained. (3) Overall, our method enables the accumulation of many mutations in a single round of experimentation, whereas directed evolution is largely limited to one mutation at a time. For example, on the CreiLOV dataset, the generated sequences with the highest fitness had on average 66 mutations from the parent reference sequence from which the original dataset was generated, which would not be achievable with directed evolution. We also did not directly compare our method to supervised approaches in smaller design spaces, such as 4-site combinatorial libraries (Yang et al., 2025), as we focus here on design in larger design spaces, where existing methods are lacking. Overall, traversing large swaths of sequence space will be important for faster engineering and enabling improvements to fitness that would normally be slow with directed evolution.
@@ A.3.1 Diffusion over continuous space
Discrete diffusion models (Austin et al., 2023; Campbell et al., 2022; Lou et al., 2024) generate data in discrete spaces by reversing a predefined forward Markov process. Specifically, a family of distributions $p_{t}$ evolves according to the Markov chain
@@
Both D3PM (Austin et al., 2023) and UDLM (Schiff et al., 2024) implement a uniform transition matrix $\bm{Q}_{t}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-\bm{I}$. When $T\to\infty$, the probability distribution $p_{T}$ converges to a uniform distribution.
@@ A.4 Steering methods
Table A3: Hyperparameters used to tune the guidance/steering process. The bolded parameter was chosen as the ideal parameter for the iterative “Bayesian optimization” experiment (Fig. 6). Larger guidance parameter implements stronger guidance strength.
@@
Continuous CG $10241/\beta=64,128,256,512,$ 1024
Discrete CG $39.06251/\beta=1,2.5,6.25,\textbf{15.625},$ 39.0625
@@
Discrete DAPS $2561/\beta=16,32,64,\textbf{128},$ 256
@@
Continuous NOS $\lambda=1,10,100,200,500$
$\eta=2$
$K=10$
@@ A.4.1 Classifier guidance
Classifier guidance (Song et al., 2020) is a technique used to steer samples generated by diffusion models toward desired attributes. The primary goal is to sample from a conditional distribution $p({\mathbf{x}}|{\mathbf{y}})$, where ${\mathbf{y}}$ is a guiding signal of interest. In continuous space, this can be achieved by replacing the unconditional score function $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ at time $t$ by a conditional score function,
@@ A.4.2 Posterior sampling
Another line of guidance work (Chung et al., 2023; Mardani et al., 2024; Zhang et al., 2024) focuses on drawing samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$, where the prior distribution is modeled by a pretrained diffusion model. The conditional distribution $p({\mathbf{y}}|{\mathbf{x}})$ can either be the likelihood function of a forward model (i.e., when ${\mathbf{y}}$ is an incomplete measurement of ${\mathbf{x}}$) or an exponential distribution with respect to a reward function (i.e., $p({\mathbf{y}}|{\mathbf{x}})\propto\exp(f({\mathbf{x}})/\beta)$). The major difference between posterior sampling and classifier guidance is that it requires the reward function to be trained only on clean data ${\mathbf{x}}$.
While many works have studied posterior sampling in Euclidean space with continuous diffusion models, posterior sampling for discrete data has been less explored. We modified DAPS (Zhang et al., 2024) to enable diffusion posterior sampling in discrete-state spaces. Suppose ${\mathbf{x}}$ lies in a finite support $\mathcal{X}^{D}$, we follow the following steps:
@@ A.4.4 Direct preference optimization
For DPO with language models, we used the ranked loss function from Widatalla et al. (2024) and Stocco et al. (2024) (Eq. 21). $\pi_{\theta}$ is the policy to be updated, $\pi_{\text{ref}}$ is the original model, and $\beta$ is a tunable parameter describing the extent of drift from the reference model. The loss therefore describes the cross entropy of the ratio $\beta\log\frac{\pi_{\theta}({\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{x}})}$ and the fitness value $w$. Following Stocco et al. (2024), we calculated the ratio $r$ as the difference of the log likelihood of the sequence from the updated model minus the log likelihood of the reference model, and softmax was applied to all of the fitness values $w$. We used the default parameters from (Stocco et al., 2024) but increased the learning rate to $10^{-4}$ and used higher values of the $\beta$ parameter with finetuning for 5 epochs. Future work could explore other DPO approaches such as paired and weighted loss with other types of models.
$L_{\text{DPO}_{\text{ranked}}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{D}\sum_{k=1}^{K}\left[\beta\log\frac{\pi_{\theta}({\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{x}})}-\log\sum_{j=k}^{K}\exp\left(\beta\log\frac{\pi_{\theta}({\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{x}})}\right)\right]$ (21)
@@ A.5 Adaptive optimization algorithm
7:              Sample value function $f_{\theta}\sim\text{Uniform}(\{f_{\theta_{t,m}}\}_{m=1}^{M})$triangleright Thompson-style sampling
[…]
@@
A.6 Additional results
@@
Figure A5: Pareto boundaries demonstrate the trade-off between generating sequences with high fitness and high diversity for TrpB (A-C) and CreiLOV (D-F), showing the same experiment as Fig. 5. Error bars show standard deviation. Mean fitness and diversity were calculated based on 200 generated samples, with diversity calculated as the total number of unique and novel samples in the generated batch, out of 200. Larger circles indicate a stronger guidance strength, specified in Table. A3.

### v2

@@ A.2 Protein fitness optimization task
We studied fitness optimization across three different protein-fitness datasets, TrpB, CreiLOV, and GB1 (Table 2). TrpB is 389 residues in length, but based on available fitness data, we limited design to 15 residues: 117, 118, 119, 162, 166, 182, 183, 184, 185, 186, 227, 228, 230, 231, and 301. Namely, we combined the fitness data from 6 combinatorially complete 3-site libraries (D-I from Johnston et al. (2024)) and the 4-site library across residues 183, 184, 227, and 228. We normalized the parent fitness to 1 in each dataset and rounded all negative fitness values up to zero. The fitness here is the catalytic rate of a native reaction, the formation of tryptophan from indole and serine. To obtain a proxy fitness for all variants in the design space ($20^{15}$ possibilities) we trained an oracle inspired by the dataset splitting and model architecture used in Blalock et al. (2025). Namely, we used all of the single, double, and triple mutants in the library for training, with 10% and 20% of the quadruple mutants being used for validation and testing, respectively. Our model consists of an ensemble of 20 MLPs for TrpB, and each was trained on one-hot encodings of the designed residues for 1000 epochs.
@@
For GB1, the experimental finesses for nearly all double mutations across the entire protein were available, where fitness refers to binding affinity of a domain of the G protein. To train the oracle, we held out 10% and 20% the sequences with two mutations as a validation and test set, respectively, with remaining sequences being used for training. Our model consists of an ensemble of 10 MLPs for GB1, and each was trained on one-hot encodings of the designed residues for 50 epochs.
@@
Figure A1: Oracles trained on available labeled fitness data for TrpB, CreiLOV, and GB1 extrapolate well to higher order combinations of mutations within the design space, as measured by Pearson correlation.
@@
Our primary method for evaluation involved examining the distribution of sampled sequences and their corresponding fitness values, diversities, and novelties. The processing pipeline for generated sequences in shown in Fig. A2. In diffusion models, sequences were generated with fixed length equal to the parent length. For the language models, nearly all generated sequences had length equal to the parent sequence length. Still, sequences were aligned with the parent sequence using mafft (Katoh & Standley, 2013), and gaps were replaced with the corresponding amino acid in the parent sequence to generate complete pseudo-sequences of a fixed length. Special tokens, which occurred rarely in generation, were replaced by a random amino acid. For TrpB, residues outside of the design space of 15 residues were naively mapped to the original amino acid type in the parent sequence at the end of generation. We did not test inpainting, although this could be accomplished with masked (diffusion) language models.
@@
There are several reasons why we did not directly compare the performance of SGPO methods to existing methods used in protein engineering, such as directed evolution and MLDE. In the case of directed evolution (such as random mutagenesis): (1) It is not obvious which parent sequences to use as the starting points for directed evolution for a fair comparison. (2) It is unclear if the oracle captures the true nature of the protein fitness landscape or extrapolates well to sequences with many mutations relative to the original fitness dataset from which the oracle was trained. (3) Overall, our method enables the accumulation of many mutations in a single round of experimentation, whereas directed evolution is largely limited to one mutation at a time. For example, on the CreiLOV dataset, the generated sequences with the highest fitness had on average 66 mutations from the parent reference sequence from which the original dataset was generated, which would not be achievable with directed evolution. We also did not directly compare our method to supervised approaches in smaller design spaces, such as 4-site combinatorial libraries (Yang et al., 2025b), as we focus here on design in larger design spaces, where existing methods are lacking. Overall, traversing large swaths of sequence space will be important for faster engineering and enabling improvements to fitness that would normally be slow with directed evolution.
@@ A.3.1 Diffusion over continuous space
Discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024) generate data in discrete spaces by reversing a predefined forward Markov process. Specifically, a family of distributions $p_{t}$ evolves according to the Markov chain
@@
Both D3PM (Austin et al., 2021) and UDLM (Schiff et al., 2024) implement a uniform transition matrix $\bm{Q}_{t}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-\bm{I}$. When $T\to\infty$, the probability distribution $p_{T}$ converges to a uniform distribution.
@@ A.4 Steering methods
Table A3: Hyperparameters used to tune the guidance/steering process. The bolded parameter was chosen as the ideal parameter for the iterative “Bayesian optimization” experiment (Fig. 6). Larger guidance parameter typically implements stronger guidance strength.
@@
Continuous CG $1/\beta=64,128,256,512,1024$
Discrete CG $1/\beta=1,2.5,6.25,\textbf{15.625},39.0625$
@@
Discrete DAPS $1/\beta=16,32,64,\textbf{128},256$
@@
Continuous NOS $\lambda=0.1,1,10,100,\textbf{1000}$
$\eta=0.5,\textbf{2},5$
$K=5,\textbf{10}$
@@ A.4.1 Classifier guidance
Classifier guidance (Song et al., 2021) is a technique used to steer samples generated by diffusion models toward desired attributes. The primary goal is to sample from a conditional distribution $p({\mathbf{x}}|{\mathbf{y}})$, where $\mathbf{y}$} is a guiding signal of interest. In continuous space, this can be achieved by replacing the unconditional score function $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ at time $t$ by a conditional score function,
@@ A.4.2 Posterior sampling
Another line of guidance work (Chung et al., 2023; Mardani et al., 2024; Zhang et al., 2025) focuses on drawing samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})\propto p({\mathbf{x}})p({\mathbf{y}}|{\mathbf{x}})$, where the prior distribution is modeled by a pretrained diffusion model. The conditional distribution $p({\mathbf{y}}|{\mathbf{x}})$ can either be the likelihood function of a forward model (i.e., when $\mathbf{y}$} is an incomplete measurement of $\mathbf{x}$}) or an exponential distribution with respect to a reward function (i.e., $p({\mathbf{y}}|{\mathbf{x}})\propto\exp(f({\mathbf{x}})/\beta)$). The major difference between posterior sampling and classifier guidance is that it requires the reward function to be trained only on clean data $\mathbf{x}$}.
While many works have studied posterior sampling in Euclidean space with continuous diffusion models, posterior sampling for discrete data has been less explored. We modified DAPS (Zhang et al., 2025) to enable diffusion posterior sampling in discrete-state spaces. Suppose $\mathbf{x}$} lies in a finite support $\mathcal{X}^{D}$, we follow the following steps:
@@ A.4.4 Direct preference optimization
For DPO with language models, we used the weighted loss function from Widatalla et al. (2024) and Stocco et al. (2024) (Eq. 21). $\pi_{\theta}$ is the policy to be updated, $\pi_{\text{ref}}$ is the original model, and $\beta$ is a tunable parameter describing the extent of drift from the reference model. The loss therefore describes the cross entropy of the ratio $\beta\log\frac{\pi_{\theta}({\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{x}})}$ and the fitness value $w$. Following Stocco et al. (2024), we calculated the ratio $r$ as the difference of the log likelihood of the sequence from the updated model minus the log likelihood of the reference model, and softmax was applied to all of the fitness values $w$. We used the default parameters from (Stocco et al., 2024) and tested increasing the learning rate to $10^{-4}$ but found that generation quality broke down above the levels used in Table A3 with finetuning for 5 epochs. We also tested ranked loss with other types of models, but the performance was similar.
$L_{\text{DPO}_{\text{weighted}}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{D}\sum_{k=1}^{K}w^{k}\left[\beta\log\frac{\pi_{\theta}(x)}{\pi_{\text{ref}}(x)}-\log\sum_{j=k}^{K}\exp\left(\beta\log\frac{\pi_{\theta}(x)}{\pi_{\text{ref}}(x)}\right)\right]$ (21)
@@ A.5 Adaptive optimization algorithm
7:     Sample value function $f_{\theta}\sim\text{Uniform}(\{f_{\theta_{t,m}}\}_{m=1}^{M})$ $\triangleright$ Thompson-style sampling
8:     Sample sequence ${\mathbf{x}}_{b}\sim\text{GuidedSample}(p({\mathbf{x}}),f_{\theta},\text{GuidanceStrategy})$
@@
A.6 Latent space Bayesian optimization with
We utilized the APEXGo codebase (Torres et al., 2024), a package for training generative variational autoencoders over peptide sequences and then optimizing those sequences with latent space Bayesian optimization to maximize certain properties. We used the training code out-of-the-box to train variational autoencoders over the same MSA sequences used to train priors for discrete diffusion models in SGPO. We trained until losses plateaued, to 542, 241, and 391 epochs for TrpB, CreiLOV, and GB1, respectively. Overall, the reconstruction losses were low, and generated sequences had low perplexity and high fitness, comparable to the generative models used in SGPO. We then used this latent space and the APEXGo optimization algorithm to maximize the fitness of sequences as measured by the oracle used in SGPO benchmarking. Specifically, in our configuration, we set the number of initialization points to 100, the number of desired diverse solutions to 1, the max number of oracle calls to 800, and the batch size to 100–to mimic the iterative Bayesian optimization experiments performed in Fig. 6.
A.7 Additional results
@@
Figure A4: Additional results for GB1, corresponding to Fig. 4 and Fig. A3. The 4 positions shown correspond to the parent sequence of VDGV.
@@
Figure A6: Pareto boundaries demonstrate the trade-off between generating sequences with high fitness and high diversity for TrpB (A-C), CreiLOV (D-F), and GB1 (G-I) – showing the same experiment as Fig. 5. Error bars show standard deviation. Mean fitness and diversity were calculated based on 200 generated samples, with diversity calculated as the total number of unique and novel (previously unseen) samples in the generated batch, out of 200. Larger circles indicate a stronger guidance strength, specified in Table. A3.